# Semantic Inference-Based Deep Learning and Modeling for Earth Observation: Cognitive Semantic Augmentation Satellite Networks

### v1

@@ II-3 Advancements in Semantic Comprehension
Table I highlights the significant advancements in semantic image processing, computational semantic compression, and the fusion of semantic trajectories within the context of EO and related fields. From streamlining data searches and enhancing data integration to innovative compression methods like CORSA, these contributions emphasize the growing importance of semantic technologies in handling large-scale EO datasets. The development of semantic EO data cubes, real-time semantic trajectory processing systems, and advanced AI models such as HighDAN and SAMRS demonstrates how semantic understanding is transforming data analysis, enabling more efficient and accurate results. These innovations pave the way for improved applications across various domains, from autonomous vehicles and urban planning to medical image analysis and video surveillance. Nonetheless, challenges remain, particularly with respect to non-standard data types and limited datasets, requiring further research to optimize deep learning models and enhance generalizability across diverse environments.
III Formation of semantic communication networks
@@
The utilization of semantic communication (SemCom), which specifically prioritizes the transmission of the fundamental meaning (semantics) of information, has been increasingly popular because of its ability to create more efficient communication systems. Notwithstanding this commitment, there remains a discrepancy between theoretical principles and their practical execution. In [39], the authors present that fills the existing gap by reevaluating the core concepts of SemCom and their impact on system design. The investigation includes two primary components: a goal-oriented methodology that gives priority to metrics focused on data interpretation, and semantic operability, which considers both data sharing and its relevance. Within the field of engineering design, semantic networks play a crucial role by using knowledge graphs called nodes and links to visually represent ideas and their interconnectedness [40]. They are increasingly employed as storage places of information to support various engineering design tasks.
@@ III-1 Semantic networks evolving toward SemCom networks
Following the SemCom networks presented in [46], a semantic-aware infrastructure provides provisioning and orchestration for 6G services with stringent QoS/QoE requirements. Service providers register their services, which consist of functional instances such as delivering and monitoring of motion, and this registration is facilitated via SemCom through relevant Knowledge Bases (KBs). These services use lifelong learning activities to ensure the knowledge repositories are regularly updated. Clients seek access to these services by establishing direct connections that manage the flow of semantic information to the service instances. The reasoning plane in end nodes in advanced language maturity requires semantic depictions from the user and control planes for the succeeding multiple access command. This enhances the cognitive capacities and operational autonomy of radio nodes in their decision-making processes for networking and control.
III-2 Orientation and attributes of semantic communication
@@
The semantic attribute assesses the congruence between the recipient’s deduced understanding and the presenter’s intended message, by considering the content, specifications, and semantics, with the aim of improving the intelligence of the communication system. Coherence encompasses both the precision and lucidity with which symbols may convey their intended meanings, as well as the effective conveyance of information or a concept from its origin to its destination, without delving into particular particulars. The evaluation of the compression level attained in semantic source coding in relation to the original information is of utmost significance [49]. Through the use of this semantic source, the authors have derived the theoretical limits for both lossless and lossy compression, along with the minimum and maximum constraints on the rate-distortion function. Based on the defined research objectives and existing knowledge, it can be shown that semantic source coding [19] has the potential to achieve more significant reductions in communication overhead and redundancy cost when compared to separate source-channel coding (SSCC). The primary factor for this is its ability to collect and compress the most pertinent and concise information, and then summarize it effectively.
  • Evaluation of the level of compression achieved in semantic source coding with respect to the original information is crucial. In order to determine the theoretical boundaries for lossless and lossy compression, the writers of [49] employed this semantic source. Moreover, they established the lower and upper limits for the rate-distortion function.
  • The aim of the task-oriented component is to effectively accomplish tasks while taking into account the possible influence on the value of the information delivered. Leveraging easily accessible resources such as communication bandwidth, processing expenses, and energy consumption enhances the feasibility of attaining a shared goal within the constraints and specifications of a particular project.
  • An assessment of the effectiveness of a task or performance indicator relies on the extent to which the obtained information contributes to its achievement. System performance evaluation may be quantified by measuring the extent to which a certain objective is achieved, taking into account the allocated resources.
These previously described perspectives differ from the approach grounded on the information theory paradigm, which considers all potential transmission sequences.
@@ III-3 Semantic joint source-channel coding
Figure 2: The block diagram for semantic inference-based earth observation.
Traditional communication networks treat all data as equal and provide priority to sending it without any mistakes. This statement neglects to recognize the varying importance of the information itself. Source coding is a data compression technique used to prioritize the delivery of just the essential components. This suggests that conventional methods may be improved by giving priority to the relevancy of information. The objective of SemCom is to understand the fundamental meaning of the message. Despite its growth, the discipline has challenges such as developing models for complex data and using current communication systems. Notwithstanding the obstacles, the Joint Source-Channel Coding (JSCC) method has the capacity to enhance semantic communication in future systems like 6G.
@@
The primary goal of deep learning-based Joint Source-Channel Coding (DJSCC) [52] is to reliably attain high levels of performance, even in scenarios characterized by resource constraints and low Signal-to-Noise Ratios (SNRs). The reconstruction scheme involves the use of DJSCC, as outlined in the research by [53], which may be seen as an adapted version of an auto-encoder. Supplementary Noise Ratios (SNRs) are included in the intermediate part of the encoder, resulting in the generation of a modified version of the original auto-encoder. In contrast to traditional wireless transmission techniques, the DJSCC encoder [54] employs semantic signals, which are analog waveform sequences, instead of digital transmission. Consequently, it may operate with high efficiency even in challenging channel conditions while still achieving enough recovery. Nevertheless, the enhanced efficiency attained by modern DJSCC techniques is only evident via simulations in scholarly research, which function as a type of information dissemination via supplementary transmission for semantic networks [55]. Typically, these models make the assumption of flawless synchronization, precoding, antenna, and power amplifier conditions. Therefore, the study [56] focuses on a DJSCC platform that employs SDR technology. The evaluation of the system’s capability relies on two crucial factors, namely non-linear distortion and synchronization inconsistencies. Moreover, the authors of [57] propose an innovative method for semantic communication, drawing inspiration from the impressive adaptability shown by Vision Transformers (ViTs) in successfully addressing various issues associated with image disruptive behavior. The potential advantages outlined in [58] of incorporating Unequal Error Protection (UEP) into a semantic encoder/decoder, using the current JSCC system, are worth investigating. The objective of this approach is to preserve the relative importance of data focused on semantic tasks. Furthermore, the implementation of a JSCC system may be simplified by using Quasi-cyclic Low-Density Parity-Check (QC-LDPC) codes on an adaptable device. A prototype for semantic communications is developed by constructing a fixed-point system that is then used to transmit and receive semantic information across a channel that contains residual noise.
@@
In Fig.2, a semantic transmitter extracts and encodes meaningful information from the source, which is then modulated using light-emitting diode (LED)-based optical transmission [62]. The optical channel carries this data to a camera receiver, where it is demodulated and decoded through JSCC to recover the original semantic information. A key feature of the system is its use of feedback loops and knowledge sharing, which allow the transmitter to adapt its encoding and transmission strategies in response to channel conditions, improving robustness and efficiency. The system employs adaptive optical transmission, adjusting parameters like modulation and power levels to enhance performance under varying environmental conditions. The knowledge base in this semantic communication system plays a central role in enhancing efficiency and accuracy by focusing on transmitting meaningful information rather than raw data. The transmitter and receiver can interpret semantic representations by sharing a common knowledge base, ensuring efficient communication with reduced data transmission. The system incorporates real-time feedback, allowing the transmitter to adapt its encoding and transmission strategies based on current channel conditions, such as noise or interference.
@@ IV-2 Design of semantic effective plane
OpenAGI [66] is an open-source AGI research platform designed to solve complex, multi-step tasks by synthesizing diverse models and utilizing datasets from sources like Hugging Face, GitHub, and LangChain. It introduces the Reinforcement Learning from Task Feedback (RLTF) approach, where a Large Language Model acts as a controller to select and execute external expert models, using task feedback to refine its planning strategy. By combining LLMs with domain-adapted models, OpenAGI aims to enhance task-solving capabilities, even with smaller-scale models, advancing the potential for Artificial General Intelligence (AGI). This is achieved by analyzing existing studies on these tasks and their practical applications, as illustrated in Fig.3 which depicts the semantic effective plane of Fig.1. Our proposed semantic effective plane can be achieved by the following outlines:
Figure 3: The proposed semantic effectiveness plane applying semantic processing with LLM technique
  • By leveraging knowledge graphs, OpenAGI enriches its understanding of complex tasks through a structured representation of entities and their relationships, allowing for more accurate and context-aware decision-making. This integration facilitates a deeper comprehension of the domain-specific knowledge and connections relevant to multi-step tasks.
@@
  • The hybrid mode of knowledge graphs and pretrained predictive models empowers OpenAGI to approach problems with a well-rounded perspective, integrating domain knowledge with predictive insights. As a result, OpenAGI not only refines its task-solving strategies through RLTF but also achieves a higher level of accuracy and efficiency in addressing complex, real-world challenges. OpenAGI can formulate task-oriented semantic commands. These commands are designed to be actionable and directly relevant to the task, guiding the execution of appropriate models or tools.
In modern networking architectures, different network layers serve distinct purposes, from physical transmission to data presentation, each playing a critical role in ensuring efficient communication. The semantic importance of these layers lies in their ability to abstract and compartmentalize network functions, such as error correction, flow control, routing, and data encryption. However, cross-layer inference, where decisions at one layer affect or leak into others, can introduce inefficiencies, security vulnerabilities, or performance bottlenecks. LLMs can mitigate these issues by understanding the dependencies between layers and applying intelligent decision-making strategies. By analyzing traffic patterns, latency, and data flow across layers, LLMs can propose optimizations that enhance end-to-end performance while maintaining layer independence.
@@ IV-3 Automating Earth observation with LLMs integration
V Semantic feature processing for Resource-efficient communication and interference cancellation
This section introduces a Cognitive Semantic EO system model designed to enhance satellite communication networks by integrating advanced semantic processing techniques, such as Deep Joint Source-Channel Coding (DT-JSCC) and Semantic Data Augmentation (SA). The model leverages inter-satellite links and cognitive semantic augmentation to optimize the transmission and analysis of multispectral EO data. By focusing on transmitting essential, task-relevant semantic features rather than raw data, the system reduces communication overhead and energy consumption, while improving decision-making and performance in 6G satellite communication networks. The proposed framework showcases an end-to-end semantic-augmented communication process, ensuring efficient data handling and enhanced user experience.
V-A Cognitive Semantic EO system model and problem formulation
Figure 4: The proposed semantic cognitive satellite network applying DT-JSCC and semantic data augmentation technique.
In this satellite system illustrated in Fig.4, a low Earth orbit (LEO) satellite establishes communication with a gateway located at a ground terminal directly connected to a server. Moreover, the satellite is smoothly integrated into an intricate network of inter-satellite links, allowing it to exchange real-time data with other spacecraft in orbit. Implementing an inter-satellite connection improves data transmission reliability and extends the communication period before the satellite moves outside the ground terminal’s coverage area.
The system processes multispectral satellite images captured at different time points, utilizing two satellites. The Sat 1, positioned near the ground terminal, captures the reference image at time $t_{0}$, denoted as $U(t_{0})$. This image is represented as $U(t_{0})=\{u^{t_{0}}(i,j,k)\mid 1\leq i\leq H,1\leq j\leq W,1\leq k\leq D\}$, where $H$ is the height, $W$ is the width, and $D$ is the number of spectral bands.
V-A1 Semantic data transmission
Upon cognitive processing of the reference image $U(t_{0})$, semantic information $U_{sem}(t_{0})$ is generated by the classification network $\textit{f}_{s1}$ of on-board processor on Sat 1. This semantic information can be mathematically represented as $U_{sem}(t_{0})=\textit{f}_{s1}(U(t_{0}))$, where $U_{sem}(t_{0})=\{u^{t_{0}}(i,j,k)\mid 1\leq i\leq B,1\leq j\leq C,1\leq k\leq A\}$. We define $B$ as the batch size, $C$ as the number of classes, and $A$ as the feature dimension. This information is then sent across an inter-satellite communication connection to the second satellite, which is located at a further distance and serves as a relay. The transmission and receiving procedure can be executed using the DT-JSCC approach [69], which relies on a certain quantization level to provide an error correction for channel coding. As illustrated in Fig.5 at time $t_{1}$, Sat 2 employs the semantic information $U_{sem}(t_{0})$ transmitted by Sat 1 encoding $\textit{f}_{s2}$ to examine and establish a correlation to the current image $U(t_{1})$ that Sat 2 has obtained. This operation can be achieved by on-board training for more semantic features. This ensures that Sat 2 can effectively analyze the new image in the context of the reference image received from Sat 1.
V-A2 Cognitive semantic augmentation
Sat 2 utilizes semantic information $U_{sem}(t_{0})$ to extract and acquire pertinent features from the $U(t_{1})$ image. Semantic cognitive augmentation intends to prioritize both the pertinent data for the task and the semantic feature invariant, facilitating the effective detection of important changes and correlations between the reference and new images through the Semantic Data Augmentation (SA) methodology [70]. As illustrated in Fig.5, the covariance matrix prediction network $\textit{g}_{s2}$ learns from $U_{sem}(t_{0})$ to minimize the SA training loss $\textit{L}_{SA}(\textit{f}_{s2};\textit{g}_{s2}|U_{sem}(t_{0}))$. This can be done by setting up an SA task and optimizing $\textit{g}_{s2}$ by backpropagating the SA-gradient $\theta_{\textit{g}_{s2}}$ and minimizing the cross-entropy loss of the output of DT-JSCC decoder $\textit{l}_{s2}$. This technique aligns with the objective of the system to minimize both data transmission and on-board energy consumption [71]. Processing data on the semantic feature domain $U_{sem}(t)$, as opposed to the multi-spectral pixels domain $MP(t)$, shows potential for transmitting cognitive information effectively to satellites in close proximity. We provide the semantic cognitive covariance matrix $\upsilon_{t_{1}}^{\textit{g}_{s2}}=\textit{g}_{s2}(U_{sem}(t_{0}),\theta_{\textit{g}_{s2}})$, which enhances the generalization of image change in EO data and facilitates a smooth integration into the optimization objective of the classification network. The scope of this extends beyond particular tasks linked to the development of certain neural networks. Therefore, the cognitive semantic channel can be facilitated by the above methodology of intergrating DT-JSCC and semantic augmentation technique, which expands the applications of classification tasks for 6G communication.
Definition 1.
Cognitive semantic augmentation is the enhancement of a cognitive agent’s ability to interpret and process information by leveraging semantic knowledge. It enables the agent to prioritize relevant features, maintain semantic consistency, and improve the detection of important changes and patterns, thereby enhancing decision-making and understanding in complex environments.
Figure 5: The proposed transmitter and receiver for cognitive semantic satellite networks.
V-A3 End-to-end semantic-augmented communication
As illustrated in Fig. 4, the end-to-end communication between EO Sat 1, EO Sat 2, and the User Terminal (UT) involves several advanced steps to optimize the transmission and interpretation of EO data using the concept of Cognitive Semantic Augmentation (CSA). EO Sat 1 initiates the process by collecting signals from the Earth’s surface, which could be in the form of imagery, radar, or other sensor data. Rather than transmitting raw data, Sat 1 applies DT-JSCC to compress the data into quantized semantic features, retaining only the most essential information needed for specific tasks, thus reducing communication overhead. These semantic features are transmitted over an inter-satellite link to EO Sat 2, which is enhanced with semantic processing capabilities. Upon receiving the data, EO Sat 2 conducts on-board semantic inference and augmentation processing, analyzing and further interpreting the features to extract meaningful insights, such as detecting objects, patterns, or changes over time. The processed semantic data is then transmitted to the User Terminal (UT), which allows the UT to access this refined information. Before EO Sat 2 transmits its downlink data, UT identically received the Sat1 EO data $U_{sem}(t_{0})$ to perform SA technique [70] and optimize DT-JSCC decoder $\textit{l}_{UT}$. Therefore, the classification networks $\textit{f}_{s2}$ and $\textit{l}_{UT}$ can be updated individually as semantic-enhanced DT-JSCC to facilitate the CSA downlink. As illustrated in Fig.5, we summarize the proposed end-to-end CSA satellite communication in Algorithm 1.
Algorithm 1 End-to-end CSA Satellite communication
Input: Sat1 and Sat2 EO image $U(t_{i})$: Height $H$, Width $W$, and $D$ spectral bands; Semantic information: batch size $B$, $C$, number of classes $C$, and feature dimension $A$;
Output: UT EO inference;
1:for $i=0,1,\cdots,n-1$ do
2:   Sat1 processing and transmit to Sat2 and UT: $U_{sem}(t_{i})=\textit{f}_{s1}(U(t_{i}))$.
3:  Sat2 receive $U_{sem}(t_{i})$ and processing $U_{sem}(t_{i+1})=\textit{f}_{s2}(U(t_{i+1}))$.
4:  Covariance prediction network of Sat2: Meta-learning($\textit{g}_{s2}$,$\textit{f}_{s2}$,$\textit{l}_{s2}$) to process:
5:     $\upsilon_{t_{i+1}}^{\textit{g}_{s2}}\rightarrow$Minimise $\textit{L}_{SA}(\textit{f}_{s2};\textit{g}_{s2}|U_{sem}(t_{i}))$ and Optimizing $\textit{g}_{s2}$ by backpropagating $\theta_{\textit{g}_{s2}}$.
6:   Covariance prediction network of UT: Meta-learning($\textit{g}_{UT}$,$\textit{f}_{UT}$,$\textit{l}_{UT}$) to process:
7:   $\upsilon_{t_{i+1}}^{\textit{g}_{UT}}\rightarrow$Minimise $\textit{L}_{SA}(\textit{f}_{UT};\textit{g}_{UT}|U_{sem}(t_{i}))$ and Optimizing $\textit{g}_{UT}$ by backpropagating $\theta_{\textit{g}_{UT}}$.
8:   Sat2 transmit SA $U_{sem}(t_{i+1})$ to UT $\textit{l}_{UT}$ for SA EO inference.
9:end for
Once the semantic inference is complete, the data undergoes semantic reconstruction, processing, and analysis to produce actionable insights, which are then made available to user applications through Semantic Feature Division Access (SFDMA) [72]. These user applications can use the semantically processed EO data for a variety of purposes, including inter-UE (User Equipment) interference cancellation, which improves communication efficiency and signal clarity. The SFDMA network reduces interference through approximate orthogonal transmission and introduces the Alpha-Beta-Gamma formula for semantic communications, with simulations validating its effectiveness. This overall process demonstrates how integrating semantic cognition into satellite systems can drastically improve the efficiency of EO data transmission, interpretation, and utilization, minimizing unnecessary data loads and focusing on task-relevant information.
V-A4 On-board processing efficiency
The authors presented in [73] provide a detailed comparison of several deep learning models, focusing on metrics like input size, total parameters, estimated model size, FLOPs, and the time taken for training and inference. Simpler architectures, such as CNN and ResNet-50, exhibit low computational complexity. For instance, CNN has the smallest total parameters (0.07 MB) and FLOPs (0.93 MFLOPs), making it highly efficient for training (233 seconds) and inference (7 ms). ResNet-50, while more complex, still maintains moderate efficiency, with 195.7 million parameters, a model size of 100 MB, and 11.78 MFLOPs, resulting in faster inference times (0.83 s) compared to more complex models.

### v3

@@ II-3 Advancements in Semantic Comprehension
Table I highlights the significant advancements in semantic image processing, computational semantic compression, and the fusion of semantic trajectories within the context of EO and related fields. From streamlining data searches and enhancing data integration to innovative compression methods like CORSA, these contributions emphasize the growing importance of semantic technologies in handling large-scale EO datasets. The development of semantic EO data cubes, real-time semantic trajectory processing systems, and advanced AI models such as HighDAN and SAMRS demonstrates how semantic understanding is transforming data analysis, enabling more efficient and accurate results. These innovations pave the way for improved applications across various domains, including autonomous vehicles, urban planning, and environmental monitoring. Nonetheless, challenges remain, particularly regarding non-standard data types and limited datasets, necessitating further research to optimize deep learning models and enhance their generalizability across diverse environments.
III Establishment of networks for semantic communication
@@
Due to its capacity to establish more effective communication systems, the use of semantic communication (SemCom), which in particular emphasizes the conveyance of the information’s basic meaning (semantics), has grown in popularity. Despite this dedication, there is still a gap between the theoretical concepts and how they are really implemented. The authors of [39] close the gap by reassessing the fundamental ideas of SemCom and how they affect system design. The investigation includes two primary components: a goal-oriented methodology that gives priority to metrics focused on data interpretation [40], and semantic operability, which considers both data sharing and its relevance. Within the field of engineering design, semantic networks play a crucial role by using knowledge graphs called nodes and links to visually represent ideas and their interconnectedness [41]. They are increasingly employed as storage places of information to support various engineering design tasks.
@@ III-1 Semantic networks evolving toward SemCom networks
According to the SemCom networks outlined in [47], a semantic-aware infrastructure enables the provision and orchestration of 6G services with rigorous Quality of Service (QoS) and Quality of Experience (QoE) standards. Service providers catalog their offers, including functional instances for activities like motion delivery and monitoring. The registration procedure is facilitated by pertinent Knowledge Bases (KBs), which may incur additional costs for computation, storage, and communication. Maintaining these knowledge bases necessitates computing resources for data updating and querying, storage capacity for information retention, and communication bandwidth for transmitting semantic data between nodes and service instances. To tackle these difficulties, services use lifelong learning initiatives to maintain the currency of knowledge repositories. Clients get these services by creating direct links that regulate the flow of semantic information to the service instances. In this scenario, the reasoning layer in terminal nodes, endowed with sophisticated linguistic skills, necessitates semantic representations from both user and control layers to enable various access instructions. This integration augments the cognitive abilities and operational independence of radio nodes, hence enhancing their decision-making processes for networking and control.
III-2 Semantic communication perspective and characteristics
@@
The semantic attribute assesses the congruence between the recipient’s deduced understanding and the presenter’s intended message, by considering the content, specifications, and semantics, with the aim of improving the intelligence of the communication system. Coherence encompasses both the precision and lucidity with which symbols may convey their intended meanings, as well as the effective conveyance of information or a concept from its origin to its destination, without delving into particular aspects. The evaluation of the compression level attained in semantic source coding in relation to the original information is of utmost significance [50]. Through the use of this semantic source, the authors have derived the theoretical limits for both lossless and lossy compression, along with the minimum and maximum constraints on the rate-distortion function. Based on the defined research objectives and existing knowledge, it can be shown that semantic source coding [18] has the potential to achieve more significant reductions in communication overhead and redundancy cost when compared to separate source-channel coding (SSCC).
@@ III-3 Semantic joint source-channel coding
Figure 2: The block diagram for the proposed semantic inference-based earth observation adapting the techniques in [51].
Traditional communication networks treat all data as equal and provide priority to sending it without any mistakes. This statement neglects to recognize the varying importance of the information itself. Source coding is a data compression technique used to prioritize the delivery of just the essential components. This suggests that conventional methods may be improved by giving priority to the relevancy of information. The objective of SemCom is to understand the fundamental meaning of the message. Despite its growth, the discipline has challenges such as developing models for complex data and using current communication systems. Notwithstanding the obstacles, the JSCC method has the capacity to enhance semantic communication in future systems like 6G.
@@
The primary goal of deep learning-based Joint Source-Channel Coding (DJSCC) [54] is to reliably attain high levels of performance, even in scenarios characterized by resource constraints and low Signal-to-Noise Ratios (SNRs). The reconstruction scheme involves the use of DJSCC, as outlined in the research by [55], which may be seen as an adapted version of an auto-encoder. Supplementary noise ratios are included in the intermediate part of the encoder, resulting in the generation of a modified version of the original auto-encoder. In contrast to traditional wireless transmission techniques, the DJSCC encoder [56] employs semantic signals, which are analog waveform sequences, instead of digital transmission. Consequently, it may operate with high efficiency even in challenging channel conditions while still achieving enough recovery. Nevertheless, the enhanced efficiency attained by modern DJSCC techniques is only evident via simulations in scholarly research, which function as a type of information dissemination via supplementary transmission for semantic networks [57]. Typically, these models make the assumption of flawless synchronization, precoding, antenna, and power amplifier conditions. Therefore, the study [58] focuses on a DJSCC platform that employs software-defined radio technology. The evaluation of the system’s capability relies on two crucial factors, namely non-linear distortion and synchronization inconsistencies. Moreover, the authors of [59] propose an innovative method for semantic communication, drawing inspiration from the impressive adaptability shown by Vision Transformers (ViTs) in successfully addressing various issues associated with image disruptive behavior. The potential advantages outlined in [60] of incorporating Unequal Error Protection (UEP) into a semantic encoder/decoder, using the current JSCC system, are worth investigating. The objective of this approach is to preserve the relative importance of data focused on semantic tasks. Furthermore, the implementation of a JSCC system may be simplified by using Quasi-cyclic Low-Density Parity-Check (QC-LDPC) codes on an adaptable device. A prototype for semantic communications is developed by constructing a fixed-point system that is then used to transmit and receive semantic information across a channel that contains residual noise.
@@
The integration of semantic technologies in the fast-growing area of communication systems offers a novel method for improving data transmission efficiency. Our proposed technique presents a semantic transmitter that collects and encodes significant information directly from the source, enhancing the data preparation for transmission. This strategy utilizes the distinctive characteristics of light-emitting diodes (LEDs) in optical communication to enhance information transmission [51], especially in situations where conventional approaches may be hindered by noise and bandwidth constraints. The semantic transmitter functions by examining the content and context of the data before modulation. This facilitates the extraction of pertinent elements that enhance the display of information, hence improving decision-making for the recipient. The modulation method then converts these semantic representations into a format appropriate for LED-based optical transmission, assuring the preservation of the information’s integrity and relevance throughout the communication process.
As illustrated in Fig.2, a semantic transmitter extracts and encodes meaningful information from the source, which is then modulated using light-emitting diode (LED)-based optical transmission [51]. The optical channel carries this data to a camera receiver, where it is demodulated and decoded through JSCC to recover the original semantic information. A key feature of the system is its use of feedback loops and knowledge sharing, which allow the transmitter to adapt its encoding and transmission strategies in response to channel conditions, improving robustness and efficiency. The system employs adaptive optical transmission, adjusting parameters like modulation and power levels to enhance performance under varying environmental conditions. The knowledge base in this semantic communication system plays a central role in enhancing efficiency and accuracy by focusing on transmitting meaningful information rather than raw data. The transmitter and receiver can interpret semantic representations by sharing a common knowledge base, ensuring efficient communication with reduced data transmission. The system incorporates real-time feedback, allowing the transmitter to adapt its encoding and transmission strategies based on current channel conditions, such as noise or interference.
@@ IV-2 Design of semantic effective plane
OpenAGI [67] is an open-source Artificial General Intelligence (AGI) research platform designed to solve complex, multi-step tasks by synthesizing diverse models and utilizing datasets from sources like Hugging Face, GitHub, and LangChain. It introduces the Reinforcement Learning from Task Feedback (RLTF) approach, where a Large Language Model acts as a controller to select and execute external expert models, using task feedback to refine its planning strategy. By combining LLMs with domain-adapted models, OpenAGI aims to enhance task-solving capabilities, even with smaller-scale models, advancing the potential for AGI. This is achieved by analyzing existing studies on these tasks and their practical applications, as illustrated in Fig.3 which depicts the semantic effective plane of Fig.1. Our proposed semantic effective plane can be achieved by the following outlines:
Figure 3: The proposed semantic effectiveness plane applying semantic processing with RAG LLM technique
  • By leveraging knowledge graphs, OpenAGI enriches its understanding of complex tasks through a structured representation of entities and their relationships, allowing for more accurate and context-aware decision-making. Integrating Retrieval-Augmented Generation (RAG) [68] enhances this approach by enabling OpenAGI to pull in relevant external information from large-scale databases dynamically, supplementing its internal knowledge base with the latest domain-specific insights. This combination of knowledge graph structuring and RAG-powered information retrieval facilitates a deeper comprehension of the domain-specific knowledge and connections relevant to multi-step tasks, allowing OpenAGI to adapt to evolving contexts and deliver precise, informed responses across varied applications
@@
  • The hybrid mode of knowledge graphs and pretrained predictive models empowers OpenAGI to tackle Earth Observation challenges from a comprehensive perspective, integrating domain knowledge with predictive insights. This integration enables OpenAGI to refine its task-solving strategies through Reinforcement Learning with Task Feedback (RLTF), leading to enhanced accuracy and efficiency in addressing complex, real-world scenarios.
OpenAGI represents a significant advancement in the integration of artificial intelligence and EO technologies. By utilizing a hybrid approach that combines knowledge graphs with pretrained predictive models, OpenAGI can tackle complex environmental challenges with greater precision and relevance. For instance, OpenAGI can generate task-oriented semantic commands such as ”identify and classify land cover types in the provided satellite imagery” or ”detect and monitor changes in water bodies over time.” These actionable commands guide the execution of appropriate models or tools, ensuring that the system effectively leverages both the knowledge from the knowledge graph and the predictive capabilities of pretrained models. By applying this approach, OpenAGI can enhance tasks such as disaster response planning, resource management, and environmental monitoring, ultimately contributing to more informed decision-making in Earth Observation applications.
@@ IV-3 Automating earth observation with LLMs integration
V On-board deep learning and hardware design for resource-efficient EO systems
This section introduces the deep learning models for on-board processing are revisited based on complexity and efficiency, focusing on metrics like model size, parameters, and training/inference time. Furthermore, recent advancements in on-board processing hardware, including SRAM-based FPGAs, radiation-hardened processors, and reconfigurable MPSoCs, have enabled more resilient and adaptable computing for space missions. These systems enhance task performance and mitigate radiation effects, ensuring reliable operation even in harsh space environments. Collectively, these technologies support efficient data handling and communication, aligning with modern satellite objectives and expanding possibilities for advanced space exploration applications.
V-A Revisit on the effectiveness of on-board EO deep learning